You bought a $1,000 par IBM callable bond in September 2019 that is callable at 103 in 2023, 102 in 2024, 101 in 2025 and at par thereafter. It has a coupon rate of 3.0%/year. Excluding interest, how much would you receive if IBM called the bond in: a)2023 b)2024 c)2025 c)2026 e)When is IBM likely to call the bond?
Answer #1

Answer:

Given that:

The bond is callable at 103 in 2023, 102 in 2024, 101 in 2025 and at par thereafter.

Bond par value = $1,000

Excluding interest amount you would receive :

(a) in 2023 =1000 * 103% = $1,030

(b) in 2024 = 1000 * 102% = $1,020

(c) in 2025 = 1000 * 101% = $1,010

(d) in 2026 = $1,000

Hence:

Excluding interest amount you would receive :

(a) in 2023 = $1,030

(b) in 2024 = $1,020

(c) in 2025 = $1,010

(d) in 2026 = $1,000

Answer (e): IBM is likely to call the Bond when interest rate moves in favorable direction for them and it is more beneficial for them to call the bond and borrow at more favorable rate.